OpenStudy (anonymous):

Find the dot product a = 6i + 5j, b = -5i + 4j

OpenStudy (caozeyuan):

well, A*B=a1b1+a2B2 where A=(a1,a2) B=(b1,b2)

OpenStudy (astrophysics):

\[a \cdot b = (6 \times -5)i+(5 \times 4) j\]

OpenStudy (astrophysics):

It's the same thing as matrix multiplication if you've done that

OpenStudy (caozeyuan):

@Astrophysics isn't dot suppose to yiled scalar not vector?

OpenStudy (astrophysics):

Yes, it is a scalar